Quad Enterprises is considering a new three-year expansion project that requires an initial fixed asset investment of $2.31 million. The fixed asset will be depreciated straight-line to zero over its three-year tax life. The project is estimated to generate $1,725,000 in annual sales, with costs of $635,000. The project requires an initial investment in net working capital of $280,000, and the fixed asset will have a market value of $225,000 at the end of the project.

a. If the tax rate is 23 percent, what is the project’s Year 0 net cash flow? Year 1? Year 2? Year 3? (Do not round intermediate calculations and enter your answers in dollars, not millions of dollars, e.g., 1,234,567. A negative answer should be indicated by a minus sign.)
b.

If the required return is 11 percent, what is the project's NPV? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

a). Depreciation = Cost of the Project / Useful life years = $2,310,000 / 3 = $770,000

Cash flow at year 0 = Cost of Project + Investment in NWC = $2,310,000 + $280,000 = $2,590,000

Annual Cash Flow(Year 1 - 3) = [{Sales - Costs} * (1 - t)] + [Depreciation * t]

= [{$1,725,000 - $635,000) * (1 - 0.23)] + [$770,000 * 0.23]

= $839,300 + $177,100 = $1,016,400

Additional Cash Flow at year 3 = After-tax Salvage Value + Recovery of NWC

= [$225,000 * (1 - 0.23)] + $280,000 = $173,250 + $280,000 = $453,250

b). NPV = PV of Cash Inflows - PV of Cash Outflows

= [$1,016,400 * {1 - (1 + 0.11)-3} / 0.11] + [$453,250 / (1 + 0.11)3] - $2,590,000

= [$1,016,400 * 2.4437] + [$453,250 / 1.3676] - $2,590,000

= $2,483,791.64 + $331,412.49 - $2,590,000 = $225,204.13